# The Role of Accuracy and Validation Effectiveness in Conversational Business Analytics

### v2

@@ 3.3 Validation Effectiveness
When delegating information production to human experts, there is a challenge of ”hidden action,” where the actions of experts are not fully observable or assessable, potentially leading to misalignment of interests [27]. Similarly, when delegated to AI, the issue shifts also to the transparency and reliability of outputs. While human experts may act in self-interest, leading to agency costs through false information or requiring additional incentives, AI systems present the challenge of output opacity. End users cannot evaluate the correctness of SQL queries and resulting KPIs until after implementation, risking decisions based on false information with economic consequences. Ensuring the reliability of AI-generated information is crucial, particularly in high-stakes decision-making scenarios where false information can result in significant negative business outcomes. The requirement for reliability is further intensified by the inherent tendency of LLMs to produce hallucinations [43, 63], resulting in fabricated or erroneous SQL queries that may appear credible. Consequently, robust techniques are essential to ensure the validation of these generated SQL queries.
@@
The techniques for validating automatically generated SQL queries fall within the broader field of explainable AI [15, 67, 49, 37, 24]. Since the advent of expert systems, researchers have emphasized the pivotal role of explanations in enhancing the interpretability of AI-generated results [83, 55]. In recent years, the growing demand for transparency in machine learning has prompted significant advances in the field of explainable AI, leading to an expansion in both the scope and sophistication of explanation techniques. The primary objective of explainable AI is to provide explanations that elucidate the underlying processes behind AI outputs, thereby facilitating tasks such as debugging, regulatory compliance, the establishment of user trust, and the effective utilization of AI.
@@
In the specific context of Text-to-SQL, validation necessitates the provision of explanations that help to comprehend and verify AI-generated SQL queries. These explanations must detail the structure of the queries, including the involved tables, fields, and operations (e.g., filtering, grouping, or joining). Such detailed information enables users to critically evaluate the logical coherence of the query and its alignment with their intended objectives, thereby supporting an independent validation process. Effective validation ensures that users can reliably differentiate between correct and incorrect SQL queries. However, validation is ineffective if incorrect SQL queries are misclassified as correct, or correct ones are misclassified as incorrect.
@@
Validation in Text-to-SQL extend beyond mere explanation, often encompassing also techniques for error correction. These techniques reflect the interactive nature of CBA and aim to support users in both identifying and rectifying errors in SQL queries. In this study, the interpretability of AI outputs and their correction are treated separately, with a focus on validation effectiveness as the primary metric. Validation serves as the foundation for reliable decision-making, as correct validation is essential for subsequent error correction.
@@
Current literature identifies three different techniques for validating and explaining SQL queries (see [60] for an overview). Decomposition techniques (see Figure 5) break SQL queries into components such as tables, fields, and operations, presenting intermediate results based on sample data to trace each element’s contribution to the final output [57]. Visualization techniques use graphical representations to map relationships among query components, enhancing clarity and navigation in complex query structures [39, 48, 53]. Dialogue techniques employ LLMs to provide natural language explanations, enabling users to interactively explore and refine queries by addressing specific components or the underlying data model [75, 20, 36, 26, 84, 90].
@@
Empirical studies on validation techniques for Text-to-SQL reveal varying levels of effectiveness in supporting SQL comprehension and error correction. Ning et al., through a user study, observed that decomposition, visualization, and dialogue techniques achieved an effectiveness rate of approximately 56% [60]. In contrast, Tian et al. demonstrated that their technique, which combines aspects of decomposition and dialogue techniques, significantly outperformed existing techniques, achieving a validation and error correction effectiveness of approximately 85% [75]. These findings underscore substantial progress in user-based validation and error correction, showcasing their transformative potential to enhance the reliability of AI-generated SQL queries.
@@
Tian et al. have shown that their technique consistently delivers effective results regardless of the end users’ technical competency [75]. However, the boundaries of user-based validation in complex business contexts, such as in data warehouses with hundreds of interrelated tables and fields, require further research. The expertise and experience of the end users play a central role in their understanding of the software code and, consequently, in their ability to evaluate the correctness of the generated information. The explanatory information in this context differs significantly from explainable AI used to interpret machine learning models. Explanations in machine learning are typically tied directly to business language, such as: "The model predicts that the customer will have poor creditworthiness because the customer’s features exhibit these values." In contrast, the explanatory information in Text-to-SQL has a distinct technical focus, aiming to enhance the user’s understanding of how the underlying SQL query functions.
@@
Furthermore, the complexity of SQL queries, particularly those involving numerous subqueries and joins, combined with the cognitive load they impose on users, affects their ability to comprehend the software code (see, for example, [52], which references developers rather than end users). This cognitive load is further exacerbated by the architecture of the underlying data model [8], especially in cases where data models include extensive dependencies and abstract schema structures that pose significant challenges for end users.
@@
Additionally, end users typically perform validation only occasionally when generating new SQL queries to meet a specific information need. Consequently, they are unable to develop long-term knowledge of SQL or gain a deeper understanding of the underlying data model. This issue has also been observed in traditional self-service analytics [82], where users were unable to build experiential knowledge of the tools and data models employed for individual information generation.
@@ 4.1 Basic Assumptions
This study introduces theoretical models based on rational choice theory, with an emphasis on expected utility theory [81, 78], to evaluate the effective use of CBA. The analysis focuses on the example of Text-to-SQL. Two primary influencing factors are considered:
@@
It is assumed that the end user knows the levels of these factors determining the performance of the delegation of insight-generating tasks to AI at hand and can judge the suitability accordingly. Delegation to a data engineer is considered as an alternative to using AI for insight generation.
@@
In an ideal scenario, both accuracy and validation effectiveness would be perfect, ensuring the generation of reliable and actionable information. However, these factors are inherently probabilistic and subject to imperfections. The models examine how the interaction between accuracy and validation effectiveness influences the effectiveness of Text-to-SQL and its potential advantages over delegation to a human expert. Specifically, the models evaluate how varying levels of these factors affect the relative advantage of delegating tasks to AI, with or without validation support. By exploring this interplay, the models aim to identify the conditions under which Text-to-SQL can serve as a viable alternative to human expertise, providing a structured framework for its adoption in CBA.
@@
The models assume a risk-neutral perspective, where the end user has a linear utility function. This simplification allows for a focus on average outcomes (expected values) while ignoring the variability in potential gains and losses. By assuming linearity, the models quantify all values in monetary units.
@@
The analysis begins with the end user’s need for information that cannot be provided by standard reporting. The current value of a KPI is critical for making informed business decisions. While users may have access to data marts, the necessary data resides in data warehouses, requiring an SQL query for extraction. This query retrieves and prepares the data, ultimately generating the required KPI. A correctly formulated SQL query produces a valid KPI, enabling optimal decisions and yielding a net profit of $+1$. Conversely, an erroneous query generates a misleading KPI, resulting in suboptimal decision and a financial loss of $-1$. The analysis assumes that the data warehouse contains high-quality, undistorted data, making the correctness of the SQL query the sole determinant of information quality. Despite possessing expertise in business processes, the end user lacks technical skills required to apply traditional self-service analytics effectively.
@@
Currently, no data engineer is available to create the KPI, and such a resource will only be accessible at a later time. The delayed delivery of the KPI results in a reduced (net) profit $v$ (where $10<v<$ 1). This profit is lower than the achievable maximum of $+1$, reflecting the adverse effects of delayed action. In extreme cases, $v$ may even approach zero, rendering the KPI almost valueless. Additional factors, such as the effort required to align requirements and potential motivational challenges that could negatively impact collaboration between end user and data engineer, are not considered. Consequently, $v$ serves as a measure of the urgency of KPI procurement, with lower values of $v$ indicating greater urgency.
@@
The process of KPI generation can be delegated to AI equipped with a natural language interface powered by LLMs. AI offers two levels of assistance: partial support (PS) and full support (FS). Under PS, the AI translates the natural language request into SQL code, which is executed on the data warehouse to calculate the KPI. The KPI is then used to inform decision-making, resulting in a business impact. FS includes an additional validation step to ensure the correctness of the SQL query. During this step, the end user is provided with explanations of the generated SQL query, which are designed to ensure transparency and enable independent validation of its correctness. If the SQL query is deemed correct, the process concludes with confirmation, allowing to act on the KPI. If the query is identified as erroneous, it is rejected, and no action is taken, avoiding both gain and loss.
@@
In contrast to the dynamic, iterative nature of CBA, this study adopts a static perspective in which KPI generation and validation occur in a single execution cycle. This idealized approach isolates the roles of information generation and validation, enabling a focused analysis of their individual impacts on the effectiveness of Text-to-SQL. The static perspective is also chosen, as iterative processes for information generation and validation within CBA are still under development.
@@
It is further assumed that the end user incurs no direct costs (e.g., for data storage, processing, or the use of the LLM) when utilizing Text-to-SQL. This aligns with the prevailing practice in traditional self-service analytics, where no fees are charged for individual instances of information generation.
@@
Since traditional self-service analytics is not feasible due to a lack of expertise, the end user evaluates whether the KPI should be generated via delegation to AI (either PS or FS), or whether it is more advantageous to wait for the data engineer to receive the KPI at a later time. The end user selects the option with the highest benefit. When CBA (either PS or FS) offers a higher expected value than the profit through the data engineer, it is preferred. Conversely, if the expected value from CBA is lower, human delegation is preferred. In such cases, CBA is deemed ineffective, as it fails to provide timely and decision-relevant information.
@@
The following table provides the mathematical symbols used in the following and their explanation.
@@
Table 2: Overview of the mathematical symbols used and their meaning.
@@ 4.2 Partial Support (PS)
The success of this approach depends primarily on the accuracy of the AI in translating natural language requests into correct SQL queries. As outlined earlier, current advancements such as fine-tuning combined with prompt design are employed to enhance the AI’s information generation capabilities. Given the probabilistic nature of LLMs, the generation of SQL queries is treated as a discrete random variable. The probability of generating a correct SQL query, denoted as $\alpha$ (where $10<\alpha<$ 1), corresponds to execution accuracy. This probability is determined by the ratio of successfully generated SQL queries to the total number of natural language requests of a similar type.
The expected value $E_{PS}(\alpha)$ under PS is calculated as the weighted sum of the possible outcomes. A correct SQL query yields a profit of $+1$, while an incorrect SQL query results in a loss of $-1$. These outcomes are weighted by the probabilities of success ($\alpha$) and failure ($1-\alpha$), respectively. The expected value is therefore expressed as:
@@
The threshold (1) represents the accuracy that muss be exceeded for PS to be more advantageous than relying on a data engineer. When this condition is met ($\alpha>\alpha^{*}_{PS}$), PS is viable and enables effective KPI generation without relying on human expertise. This is particularly beneficial since traditional self-service analytics is constrained by the end user’s limited technical skills. Conversely, if this condition is not satisfied ($\alpha\leq\alpha^{*}_{PS}$), the expected value from PS falls below the profit achievable through delegation to a data engineer, rendering PS unsuitable for time-sensitive, decision-critical business analytics.
@@ 4.3 Full Support (FS)
In the second phase, the AI-generated SQL query is reviewed by the user to determine whether a correct KPI has been generated. Techniques such as decomposition, visualization, and dialog-based interaction are currently being explored to support users in comprehending the SQL query. This also aims to achieve transparency with regard to the generation of information. However, user-based validation also entails uncertainty. Both the user’s comprehension and the complexity of the SQL query and its underlying data model significantly influence the success of validation – specifically, the ability to correctly identify valid and erroneous SQL queries. Additionally, an inherent uncertainty arises in dialog-based interactions because AI-generated explanations are based on probabilities rather than deterministic rules.
To address this uncertainty, validation effectiveness is modeled as a random variable denoted as $\beta$ (where $10<\beta<$ 1). This probability reflects the likelihood that a user can correctly assess the SQL query based on the support provided. It is defined as the ratio of successful validations to the total number of validation attempts. By adopting this probabilistic framework, the model accounts for the inherent uncertainties associated with validation processes.
The end user only takes action if an SQL query is classified as correct. The probability of correctly classifying a correct SQL query is $\alpha\cdot\beta$, resulting in a gain of $+1$. Conversely, the probability of incorrectly classifying an incorrect SQL query as correct is $(1-\alpha)\cdot(1-\beta)$, resulting in a loss of $-1$. However, if a correct SQL query is mistakenly classified as incorrect, or an incorrect SQL query is correctly classified as incorrect, no action is taken, and neither a gain nor a loss occurs.
This, the expected value $E_{FS}(\alpha,\beta)$ for FS is given by:
$E_{FS}(\alpha,\beta)=\alpha\beta\cdot(+1)+(1-\alpha)(1-\beta)\cdot(-1)=\alpha+\beta-1.$
The expected value $E_{FS}(\alpha,\beta)$ is influenced by two key factors: accuracy and validation effectiveness. Higher values of $\alpha$ indicate improved accuracy in generating correct SQL queries, while higher values of $\beta$ represent enhanced effectiveness in validating the correctness of those queries. The combined improvement in these factors increases the expected value of FS, making it more likely to yield reliable and actionable KPI. This relationship highlights the complementary roles of accuracy and validation effectiveness in ensuring the success of FS in CBA.

### v3

@@ 3.3 Validation Effectiveness
When delegating information production to human experts, there is a challenge of ”hidden actions” where the actions of experts are not fully observable or assessable, potentially leading to misalignment of interests [27]. Similarly, when delegated to AI, the issue shifts to the transparency and reliability of outputs. While human experts may act in self interest, leading to agency costs through false information or additional incentives, AI systems present the challenge of output opacity. End users cannot evaluate the correctness of SQL queries and the resulting KPIs until after the implementation, risking decisions based on false information with economic consequences. Ensuring the reliability of AI-generated information is crucial, particularly in high-stakes decision-making scenarios where false information can result in significant negative business outcomes. The requirement for reliability is further intensified by the inherent tendency of LLMs to produce hallucinations [43, 63], resulting in fabricated or erroneous SQL queries that may appear credible. Consequently, robust techniques are essential for validating the generated SQL queries.
@@
Techniques for validating automatically generated SQL queries fall within the broader field of explainable AI [15, 67, 49, 37, 24]. Since the advent of expert systems, researchers have emphasized the pivotal role of explanations in enhancing the interpretability of AI-generated results [83, 55]. In recent years, the growing demand for transparency in machine learning has prompted significant advances in the field of explainable AI, leading to an expansion of both the scope and sophistication of explanation techniques. The primary objective of explainable AI is to provide explanations that elucidate the underlying processes behind AI outputs, thereby facilitating tasks such as debugging, regulatory compliance, the establishment of user trust, and the effective utilization of AI.
@@
In the specific context of Text-to-SQL, validation necessitates the provision of explanations that help to comprehend and verify AI-generated SQL queries. These explanations must detail the structure of the queries, including the tables, fields, and operations (e.g., filtering, grouping, or joining) involved. Such detailed information enables users to critically evaluate the logical coherence of a query and its alignment with their intended objectives, thereby supporting an independent validation process. Effective validation ensures that users can reliably differentiate between the correct and incorrect SQL queries. However, validation is ineffective if incorrect SQL queries are misclassified as correct, or if correct queries are misclassified as incorrect.
@@
Validation in Text-to-SQL extends beyond mere explanation and often encompasses techniques for error correction. These techniques reflect the interactive nature of CBA, and aim to support users in identifying and rectifying errors in SQL queries. In this study, the interpretability of the AI outputs and their correction are treated separately, with a focus on validation effectiveness as the primary metric. Validation serves as the foundation for reliable decision-making, as correct validation is essential for subsequent error correction.
@@
The current literature identifies three techniques for validating and explaining SQL queries (see [60] for an overview). Decomposition techniques (see Fig. 5) break SQL queries into components such as tables, fields, and operations, presenting intermediate results based on sample data to trace each element’s contribution to the final output [57]. Visualization techniques use graphical representations to map relationships among query components, thereby enhancing clarity and navigation in complex query structures [39, 48, 53]. Dialogue techniques employ LLMs to provide natural language explanations, enabling users to interactively explore and refine queries by addressing specific components or underlying data model [75, 20, 36, 26, 84, 90].
@@
Empirical studies on validation techniques for Text-to-SQL reveal varying levels of effectiveness in supporting SQL comprehension and error correction. In a user study, Ning et al. observed that decomposition, visualization, and dialogue techniques achieved an average effectiveness rate of approximately 56% in detecting and correcting errors [60]. In contrast, Tian et al. demonstrated that their technique, which combines elements of decomposition and dialogue techniques, significantly outperformed existing techniques, achieving an average effectiveness in validation and error correction of approximately 85% [75]. These findings underscore the substantial progress in user-based validation and error correction, showcasing their transformative potential to enhance the reliability of AI-generated SQL queries.
@@
Tian et al. showed that their technique consistently delivers effective results regardless of the technical competency of end users [75]. However, the boundaries of user-based validation in complex business contexts, such as data warehouses with hundreds of interrelated tables and fields, require further research. Adding to the complexity, the explanatory information in Text-to-SQL differs significantly from the explainable AI approaches commonly used to interpret machine learning models. Explanations in machine learning are typically closely tied to business language, such as: "The model predicts that the customer will have poor creditworthiness because the customer’s features exhibit these values." In contrast, the explanatory information in Text-to-SQL has a distinctly technical focus, aiming to improve the user’s understanding of how the underlying SQL query operates. The expertise and experience of the end users play a critical role in their ability to comprehend the software code and, consequently, evaluate the correctness of the generated information.
@@
Furthermore, the complexity of SQL queries, particularly those involving numerous subqueries and joins, combined with the cognitive load they impose on users, affects their ability to comprehend software code (see, for example, [52], which references developers rather than end users). This cognitive load is further exacerbated by the architecture of the underlying data model [8], especially in cases where data models include extensive dependencies and abstract schema structures that pose significant challenges for end users.
@@
Additionally, end users typically perform validation only occasionally when generating new SQL queries to meet a new information need. Consequently, they are unable to develop long-term knowledge of SQL or to gain a deeper understanding of the underlying data model. This challenge has also been noted in traditional self-service analytics [82], where users face difficulties in building experiential knowledge of the tools and data models employed for individual information generation.
@@ 4.1 Basic Assumptions
This study introduces theoretical models grounded in rational choice theory, with a focus on expected utility theory [81, 78], to evaluate the effective use of CBA. The examination centers on the example of Text-to-SQL. Two primary influencing factors are considered in this study:
@@
It is assumed that the end user possesses knowledge of the levels of these factors, which influence how effectively AI performs in handling insight-generation tasks, and is able to assess their suitability accordingly. Delegation to a data engineer is considered as an alternative to using AI for insight generation.
@@
In an ideal scenario, both the accuracy and validation effectiveness would be perfect, ensuring the generation of reliable and actionable information. However, these factors are inherently probabilistic and are subject to imperfections. The models examine how the interaction between accuracy and validation effectiveness influences the effectiveness of Text-to-SQL and its potential advantages over delegation to a human expert. Specifically, the models evaluate how varying levels of these factors affect the relative advantage of delegating tasks to AI, with or without validation support. By exploring this interplay, the models aim to identify the conditions under which Text-to-SQL can serve as a viable alternative to human expertise, providing a structured framework for its adoption in CBA.
@@
These models assume a risk-neutral perspective, where the end user has a linear utility function. This simplification allows for a focus on the average outcomes (expected values), while ignoring the variability in potential gains and losses. By assuming linearity, the models quantify all the values in monetary units.
@@
The analysis begins with the end user’s need for information that cannot be provided through standard reporting. The current value of a KPI is critical for informed business decisions. While the end user has access to data marts, the necessary data reside in data warehouses, requiring an SQL query for extraction. This query retrieves data from the data warehouse, prepares it, and ultimately generates the required KPI. A correctly formulated SQL query produces a valid KPI, enabling optimal decisions and yielding a net profit of $+1$. Conversely, an erroneous query generates a misleading KPI, resulting in a suboptimal decision and a financial loss of $-1$. The analysis assumes that the data warehouse contains high-quality, undistorted data, making the correctness of the SQL query the sole determinant of the information quality. Despite possessing expertise in business processes, the end user lacks the technical skills required to effectively apply traditional self-service analytics.
@@
Currently, no data engineer is available to create the KPI, and such a resource will only be accessible later. The delayed delivery of the KPI results in a reduced (net) profit $v$ (where $10<v<$ 1). This profit is lower than the achievable maximum of $+1$, reflecting the adverse effects of the delayed action. In extreme cases, $v$ may even approach zero, rendering KPI almost valueless. Additional factors, such as the effort required to align requirements and potential motivational challenges that could negatively impact collaboration between end user and data engineer, are not considered. Consequently, $v$ serves as a measure of the urgency of KPI procurement, with lower values of $v$ indicating greater urgency.
@@
The KPI generation process can be delegated to an AI equipped with a natural language interface powered by LLMs. AI offers two levels of assistance: partial support (PS) and full support (FS). Under PS, the AI translates the natural language request into SQL code, which is executed on the data warehouse to calculate the KPI. The KPI is then used to inform decision-making, resulting in a business impact. FS includes an additional validation step to ensure the correctness of the SQL query. During this step, the end user is provided with explanations of the generated SQL query, which is designed to ensure transparency and enable the independent validation of its correctness. If the SQL query is deemed correct, the process concludes with confirmation, allowing to act on the KPI. If the query is identified as erroneous, it is rejected and no action is taken, avoiding both gain and loss.
@@
In contrast to the dynamic, iterative nature of CBA, this study adopts a static perspective in which KPI generation and validation occur in a single execution cycle. This idealized approach isolates the roles of information generation and validation, enabling a focused analysis of their individual effects on the effectiveness of Text-to-SQL. The static perspective was also chosen, as iterative processes for information generation and validation within CBA are still under development.
@@
It is further assumed that the end user incurs no direct costs (e.g., for data storage, processing, or the use of LLM) when utilizing Text-to-SQL. This aligns with the prevailing practice in traditional self-service analytics, in which no fees are charged for individual instances of information generation.
@@
Because traditional self-service analytics is not feasible owing to a lack of expertise, the end user evaluates whether the KPI should be generated via delegation to AI (either PS or FS), or whether it is more advantageous to wait for the data engineer to receive the KPI at a later time. The end user selects the option with the highest benefit. When CBA (either PS or FS) offers a higher expected value than the profit gained through the data engineer, it is preferred. Conversely, if the expected value from CBA is lower, human delegation is preferred. In such cases, CBA is deemed ineffective, as it fails to provide timely and decision-relevant information.
@@
Table 2 provides the mathematical symbols used in the following and their corresponding explanations.
Table 2: Mathematical symbols used and their explanations.
@@ 4.2 Partial Support (PS)
The success of PS depends primarily on the accuracy of AI in translating natural language requests into correct SQL queries. As outlined earlier, current advancements, such as fine-tuning combined with prompt design, are employed to enhance AI’s information generation capabilities. Given the probabilistic nature of LLMs, the generation of SQL queries is treated as a discrete random variable. The probability of generating a correct SQL query, denoted as $\alpha$ (where $10<\alpha<$ 1), corresponds to execution accuracy. This probability is determined by the ratio of successfully generated SQL queries to the total number of natural language requests of a similar type.
The expected value $E_{PS}(\alpha)$ under PS is calculated as the weighted sum of possible outcomes. A correct SQL query yields a profit of $+1$, whereas an incorrect SQL query results in a loss of $-1$. These outcomes are weighted by the probabilities of success ($\alpha$) and failure ($1-\alpha$), respectively. Therefore, the expected value is expressed as follows:
@@
Condition (1) defines the threshold of accuracy that must be exceeded for PS to be more advantageous than relying on a data engineer. When this condition is met ($\alpha>\alpha^{*}_{PS}$), PS is viable and enables effective KPI generation without relying on human expertise. This is particularly beneficial since traditional self-service analytics is constrained by the end user’s limited technical skills. Conversely, if this condition is not satisfied ($\alpha\leq\alpha^{*}_{PS}$), the expected value from PS falls below the profit achievable through delegation to a data engineer, rendering PS unsuitable for time-sensitive, decision-critical business analytics.
@@ 4.3 Full Support (FS)
In the second phase, the AI-generated SQL query is reviewed by the end user to determine whether a correct KPI has been generated, ensuring transparency about the functionality of the SQL query. Techniques, such as decomposition, visualization, and dialogue-based interaction, are currently being explored to support users in comprehending SQL queries. However, user-based validation entails uncertainty. Both the user’s comprehension and the complexity of the SQL query and its underlying data model significantly influence the success of validation. Additionally, inherent uncertainty arises in dialogue-based interactions because AI-generated explanations are based on probabilities rather than deterministic rules.
To address this uncertainty, the validation effectiveness is modeled as a random variable denoted as $\beta$ (where $10<\beta<$ 1). This probability reflects the likelihood that a user can correctly assess the SQL query based on the support provided. It is defined as the ratio of successful validations to the total number of validation attempts. By adopting this probabilistic framework, the model accounts for the inherent uncertainties associated with validation processes.
The end user takes action only if an SQL query is classified as correct. The probability of correctly classifying a correct SQL query is $\alpha\cdot\beta$, resulting in a gain of $+1$. Conversely, the probability of incorrectly classifying an incorrect SQL query as correct is $(1-\alpha)\cdot(1-\beta)$, which results in a loss of $-1$. However, if a correct SQL query is mistakenly classified as incorrect or an incorrect SQL query is correctly classified as incorrect, no action is taken, and neither a gain nor a loss occurs.
Thus, the end user’s expected value $E_{FS}(\alpha,\beta)$ when FS is provided is given by
$E_{FS}(\alpha,\beta)$ $=\alpha\beta\cdot(+1)+(1-\alpha)(1-\beta)\cdot(-1)$
$=\alpha+\beta-1.$
The expected value $E_{FS}(\alpha,\beta)$ is influenced by two key factors: accuracy and validation effectiveness. Higher values of $\alpha$ indicate improved accuracy in generating correct SQL queries, whereas higher values of $\beta$ represent enhanced effectiveness in validating the correctness of those queries. The combined improvement in these factors increases the expected value of FS, making it more likely to yield reliable and actionable KPI. This relationship highlights the complementary role of accuracy and validation effectiveness in ensuring the success of FS in CBA.